# Equilibrium Matching: Generative Modeling with Implicit Energy-Based Models

### v1

@@ 3.1 Training
To train an Equilibrium Matching model, we aim to construct an energy landscape in which the target gradient at ground-truth samples is zero. To do so, we first define a corruption scheme that provides a transition between data and noise. Let $\gamma$ be an interpolation factor sampled uniformly between 0 and $1$, let $\epsilon$ be Gaussian noise, and let $x$ be a sample from the training set. Denote $x_{\gamma}=\gamma x+(1-\gamma)\epsilon$ as an intermediate corrupted sample. Unlike $t$ in FM, our $\gamma$ is implicit and not seen by the model. Our goal is to define a target gradient at these intermediate samples $x_{\gamma}$ that matches an implicit energy landscape. Using a gradient direction pointing from noise to data, we write the Equilibrium Matching training objective as:
@@
where $c(\gamma)$ controls the gradient magnitude. The target gradient $(\epsilon-x)c(\gamma)$ has direction $(\epsilon-x)$ that points from noise to data and magnitude $c(\gamma)$ that vanishes as we get closer to the data manifold. We explicitly enforce $c(1)=0$, ensuring that the energy landscape has vanishing gradients at real samples. Together, the direction and magnitude result in a target gradient that supports an implicit energy landscape. When $c(\gamma)=1$, the EqM objective is exactly the negation of FM’s objective.
@@ 3.4 Implementation
Equilibrium Matching is simple to implement. We provide example pseudocode for training in Algorithm˜1 and sampling in Algorithm˜2. During training, we first compute an interpolated corrupted sample $x_{\gamma}$ from noise $\epsilon$, image $x$, and factor $\gamma$, then train the model $f$ to predict the target gradient $(\epsilon-x)c(\gamma)$ by minimizing a mean squared error objective. At inference time, Equilibrium Matching uses the predicted gradients to iteratively optimize via gradient descent.
@@
Algorithm 1 Equilibrium Matching Training
The loss function takes as input model $f$, noise $\epsilon$ (eps), image $x$, and an interpolation factor $\gamma$ (g), and returns the EqM loss.
def training_loss(f, eps, x, g):
xg = (1-g)*eps + g*x
target = (eps-x)*c(g)
@@
Algorithm 2 Equilibrium Matching Sampling
The sampling function takes as input pretrained model $f$, initialization $st$, step size $\eta$, and total steps $N$, and returns the generated sample.
def generate(f, st, eta, N):
xn = st
for i in range(N):
xn = xn - eta*f(xn)
return xn

### v2

@@ 3.1 Training
To train an Equilibrium Matching model, we aim to construct an energy landscape in which the target gradient at ground-truth samples is zero. To do so, we first define a corruption scheme that provides a transition between data and noise. Let $\gamma$ be an interpolation factor sampled uniformly between 0 and $1$, let $\epsilon$ be Gaussian noise, and let $x$ be a sample from the training set. Denote $x_{\gamma}=\gamma x+(1-\gamma)\epsilon$ as an intermediate corrupted sample. Unlike $t$ in FM, our $\gamma$ is implicit and not seen by the model. Our goal is to define a target gradient at these intermediate samples $x_{\gamma}$ that matches an implicit energy landscape. Using a gradient direction that descends from noise to data, we write the Equilibrium Matching training objective as:
@@
where $c(\gamma)$ controls the gradient magnitude. The target gradient $(\epsilon-x)c(\gamma)$ has direction $(\epsilon-x)$ that descends from noise to data and magnitude $c(\gamma)$ that vanishes as we get closer to the data manifold. We explicitly enforce $c(1)=0$, ensuring that the energy landscape has vanishing gradients at real samples. Together, the direction and magnitude result in a target gradient that supports an implicit energy landscape. When $c(\gamma)=1$, the EqM objective is exactly the negation of FM’s objective.
@@ 3.4 Implementation
Equilibrium Matching is simple to implement. We provide example pseudocode for training in Algorithm˜1 and sampling in Algorithm˜2. During training, we first compute an interpolated corrupted sample $x_{\gamma}$ from noise $\epsilon$, image $x$, and factor $\gamma$, then train the model $f$ to predict the target gradient $(\epsilon-x)c(\gamma)$ by minimizing a mean squared error objective. At inference time, Equilibrium Matching uses the predicted gradients to iteratively optimize via gradient descent. We present the pseudocode combining NAG-GD and adaptive compute.
@@
Algorithm 1 EqM Training
The loss function takes as input model $f$, image $x$, and gradient magnitude function $c$, and returns the EqM loss.
def training_loss(f, x, c):
eps = randn_like(x)
gamma = rand()
xg = (1-gamma)*eps + gamma*x
target = (eps-x) * c(gamma)
@@
Algorithm 2 Adaptive EqM Sampling (NAG)
The sampling function takes in model $f$, initial state $st$, step size $\eta$, NAG factor $\mu$, and threshold $g$, and returns the sample.
def generate(f, st, eta, mu, g):
x = st, x_last = st, grad = f(st)
while norm(grad) < g:
x_last = x
x = x - eta*grad
grad = f(x + mu*(x-x_last))
return x